# FERA: Foil Fencing Referee Assistant Using Pose-Based Multi-Label Move Recognition and Rule Reasoning

### v2

@@ 4.4 FERA-MDT
The core of FERA is FERA-MDT, an encoder-only Transformer for multi-label move and blade recognition. We describe its architecture, loss functions, training, augmentation, and hyperparameter tuning.
@@
Each input frame has 101 features (Sec. 4.3), projected to a 128-dimensional embedding. The model comprises three encoder layers, each with eight attention heads and a feed-forward dimension of 512, with dropout of 0.1. Temporal aggregation uses masked mean pooling to ensure consistent handling of variable-length sequences.
@@
Temporal structure is preserved through sinusoidal encodings with dropout:
@@
where the blade weight (0.677) was selected via Optuna to balance gradients and label counts.
@@
To improve robustness, augmentation is applied after the data is split. Methods include temporal jitter ($\Delta t\sim\{-2,\ldots,2\}$), Gaussian noise ($\epsilon\sim\mathcal{N}(0,0.05^{2})$), rotational noise ($\theta\sim U(-0.05,0.05)$), and scaling ($s\sim U(0.95,1.05)$). To mitigate imbalance, minority moves are oversampled up to 400 per class, and the majority blade class “6” is downsampled to two-thirds of the others.
@@
Hyperparameters were tuned with Optuna [1]. Across 300 trials, we explored model size, loss weighting, weight decay, and other settings. The top 10 runs were analyzed for recurring patterns, and their best-performing elements were combined into the final model.
4.5 FERA-LM
After generating move and blade predictions, FERA requires reasoning with priority rules to produce valid and explainable decisions. To address this, we introduce FERA-LM, the fourth stage of the pipeline. FERA-LM is a compact distilled language model.
Using DeepSeek-R1-Distill-Qwen-14B as the teacher model (the largest model we could run), we generated referee decisions with a structured prompt [6].
Prompt engineering for the language model.
I am an expert fencing referee. Based solely on the fencing moves and rules provided below, I will determine the correct decision. I will provide my answer in two parts as follows:
Decision: Exactly one line saying whether left or right should have priority at the end of the clip, or none
Explanation: I will provide a concise explanation of why I reached the decision in part 1, referencing moves and rules relevant to the given moves below.
I also will not include any additional commentary, train of thought text, or anything else irrelevant to the concise explanation beyond these two parts.
I will not reference rules that are not invoked in the move sequences.
At the very end of my entire answer, I will output !!! on its own line, indicating I have finished.
Rules: (Through a FAISS index [10], we pull the relevant rules from my generated index file to insert here)
Left Fencer: (moves)
Right Fencer: (moves)
My expert Referee Decision and explanation:
This prompt enforces a strict two-part response, enabling straightforward post-processing. Responses are capped at 150 tokens to ensure conciseness, and generation is forced to be deterministic for consistent outputs [21].
The distilled student model is DistilGPT-2, trained with Causal Language Modeling on data produced by the teacher. This yields a far less computationally intensive model while retaining decision quality [8]. At present, FERA-LM remains in an early stage, as its accuracy depends heavily on FERA-MDT. Future iterations will revisit distillation once move detection achieves higher F1 scores.
@@ 5 Results
We evaluate FERA in two ways: (i) an end-to-end case study demonstrating how the full pipeline produces referee decisions, and (ii) quantitative comparisons of FERA-MDT with baseline models, supplemented by ablation studies.
5.1 End-to-end Example
We first present an example clip where the referee ruled that the left fencer held priority. The prompt template is given in Section 4.5. In this case, FERA-MDT correctly identifies the actions performed by both fencers, and FERA-LM produces a concise decision and explanation that aligns with the referee’s judgment.
@@ 5.1 End-to-end Example
(a) Left fencer step forward + beat, right fencer stepping back.
(b) Left fencer lunging, right fencer counterattacking.
@@
Explanation: The left fencer initiated an attack with a step forward since the right fencer stepped backward. The left fencer then attacks with a lunge, hit (134–146) and the right fencer responds with counterattack, hit (138–148). According to the rules, the priority never shifts to the right fencer.
5.2 Evaluation metrics.
We now move from a representative clip to aggregate performance. We report four classification metrics and three calibration metrics.
@@
  • Macro-F1 $\tfrac{1}{C}\sum_{c}\text{F1}_{c}$: treats all classes equally, so rare moves (e.g., flèche) contribute as much as common ones.
  • Weighted-F1 $\sum_{c}\tfrac{n_{c}}{\sum_{j}n_{j}}\,\text{F1}_{c}$: weights each class by frequency, downplaying errors on infrequent moves.
  • Micro-F1 $\tfrac{2\sum_{c}TP_{c}}{2\sum_{c}TP_{c}+\sum_{c}FP_{c}+\sum_{c}FN_{c}}$: overall accuracy across all moves, dominated by frequent classes.
  • Hamming loss $\tfrac{1}{NC}\sum_{i=1}^{N}\sum_{c=1}^{C}\mathbf{1}[y_{i,c}\neq\hat{y}_{i,c}]$: fraction of misclassified labels (false positives + false negatives); lower is better.
For each class $c$, precision, recall, and F1-score are:
@@
ECE summarizes average calibration quality; MCE highlights worst-case miscalibration; Brier combines calibration and sharpness into a single probability-quality measure.
@@ 5.3 Move detection model comparisons
We compare FERA-MDT against three baseline architectures: a simple Transformer encoder, a BiLSTM, and a Temporal CNN (TCN).
As shown in Table 2, FERA-MDT consistently outperforms all baselines on Macro-F1 and Hamming loss, indicating stronger performance on both common and rare fencing moves. The BiLSTM performs closest to FERA-MDT, as recurrent architectures capture short-term motion dependencies well. However, given the complex multi-label structure and long-range dependencies in our setting, FERA-MDT handles them more effectively. The baseline Transformer lags due to weak regularization and lack of effective pooling, whereas FERA-MDT benefits from dropout-stabilized sinusoidal encodings, masked mean pooling, and a structured encoder stack. The TCN achieves the highest Weighted-F1, reflecting a bias toward dominant classes such as step forward and hit. For refereeing, this is undesirable since accurate recognition of rare actions is equally important.
Table 3 shows calibration results. Both the BiLSTM and baseline Transformer perform substantially worse than FERA-MDT and the TCN. Between the latter two, the TCN achieves lower ECE and MCE but a higher Hamming loss than FERA-MDT. This suggests that its apparently better calibration stems from overspecialization on frequent classes, while FERA-MDT balances rare- and common-class predictions more effectively.
Overall, FERA-MDT achieves high accuracy, robust rare-class recognition, and reasonable calibration, making it the strongest candidate for referee assistance.
@@ 5.4 Co-occurrence Analysis
Because referee decisions depend on combinations of actions, we assess whether models reproduce the ground-truth co-occurrence structure. Figures 58 compare co-occurrence mappings across architectures. The left matrix shows ground-truth labels, while the right shows model predictions; the closer the two, the better the model captures true co-occurrences. FERA-MDT produces mappings that most closely mirror the ground-truth distribution, whereas baselines tend to over-activate frequent moves and underrepresent rarer ones.
@@
These results confirm that FERA-MDT captures not only individual moves but also their distributions and interactions. We next analyze which design choices contribute most to this performance.
@@ 5.5 Ablations
To identify which components contribute most to performance, we performed ablations on FERA-MDT while keeping all other settings fixed (baselines were not ablated). We report both classification (average Macro-F1 across five folds) and calibration metrics (average ECE, MCE, and Brier score [7, 3]).
Table 4 summarizes the results. The full FERA-MDT configuration achieves the highest Macro-F1 (0.549) while maintaining strong calibration (ECE = 0.152, MCE = 0.458, Brier = 0.135). Removing per-class thresholds reduces Macro-F1 to 0.511, confirming that threshold tuning is critical. Using only the 12 normalized joints (24D) yields a Macro-F1 score of 0.534, showing that engineered features improve performance but are not solely responsible. Notably, calibration worsens substantially in the raw-joints ablation, indicating that features primarily enhance reliability with secondary accuracy gains.
Augmentation strategies provide complementary benefits. Individually, noise-only, temporal-only, and feature-specific augmentations yield Macro-F1 scores of 0.542, 0.525, and 0.537, respectively. Combining all strategies produces the best robustness and the top score (0.549). Figures 10 and 10 illustrate the instability of single augmentation: in Fold 2, the model collapses to frequent moves, while in Fold 3, it nearly matches the full model. Despite occasional strong folds, the average remains lower (0.537), underscoring the need for combined augmentation and calibration to stabilize performance.
Finally, per-class temperature scaling improves calibration without changing F1, and down-weighting blade loss slightly benefits both calibration and classification.
Overall, these ablations emphasize that FERA-MDT’s complete design—feature engineering, diverse augmentation, calibrated thresholds, and tailored loss weighting—is necessary to achieve both strong classification and reliable confidence estimates.
@@
These factors improve both accuracy and reliability, but also highlight where the system remains data- and sensing-limited, which we address next.
@@ 6 Limitations and Future Work
Our study faces several limitations, many of which suggest clear directions for future work. The primary constraint is data. With no publicly available datasets containing extensive move annotations, we relied on manual labeling, which limited model accuracy. Nevertheless, performance improved steadily as the dataset grew: with $\sim$1,000 annotated segments, F1 plateaued near 0.40, while doubling to 2,000 segments raised F1 to 0.549. This trend suggests that long-tail actions remain a bottleneck, and that expanding the dataset—especially for rare moves—will further improve accuracy.
Another limitation is that FERA relies on publicly available pose detectors, which do not support blade detection. Developing a blade-specific extractor would enable FERA-MDT to incorporate precise blade positions, thereby increasing the reliability of contact detection. At the reasoning stage, FERA-LM depends on the outputs of FERA-MDT, meaning that motion recognition errors introduce noise during both teacher generation and student distillation. Improvements in move detection should reduce this noise, leading to more accurate language-model decisions.
Generalization also remains a challenge. Our dataset consists exclusively of professional fencers, often recorded from elevated camera angles, which limits adaptability to beginners and to more varied viewpoints. Future work may incorporate 3D joint extraction to mitigate camera-angle sensitivity and extend testing to broader fencer populations.
Finally, we did not employ athlete-disjoint splits for training and validation, as we lack metadata to link clips to individual fencers. While this is partially mitigated by scale—our dataset spans roughly 130 competitors—the risk of overfitting to individuals cannot be ruled out.
Overall, these limitations outline a roadmap: larger and more diverse datasets, explicit blade detection, improved distillation from cleaner labels, and greater attention to generalization. Addressing these factors will enable both more dependable referee assistance and broader coaching applications.
@@ 7 Conclusion
We conclude by first emphasizing the motivation for developing FERA. In modern-day foil fencing, refereeing is entirely manual, with the only automated aspect being the lights, which display who hit whom. No existing automated systems have been made to assist or validate human decisions. Prior work on fencing move recognition has been limited to small subsets of footwork actions and often uses specialized sensors for extra data. Therefore, our goal was to design the first prototype referee system that operates directly from monocular video, allowing practical deployment in everyday training settings. To that end, we present (i) the first prototype AI referee for foil fencing, (ii) a new dataset with annotated moves and blade positions, and (iii) a combined pipeline that integrates pose-based recognition with rule-grounded reasoning. While FERA-MDT’s current performance (F1 = 0.549) highlights the challenge of this subject, our results demonstrate the feasibility of utilizing Transformer-based motion understanding to extract and detect actions performed by fencers under real-world scenarios. FERA could also be extended beyond refereeing. By aggregating recognized actions and blade positions, the system could recognize patterns and offer data-driven feedback to athletes and coaches. We hope this work inspires further exploration in AI-driven sports analytics.

### v3

@@ 4.4 FERA-MDT
FERA-MDT is an encoder-only Transformer for multi-label move and blade recognition.
@@
Each input frame has 101 features (Sec. 4.3), projected to a 128-dimensional embedding. The model has three encoder layers, each with eight attention heads and a feed-forward dimension of 512, with dropout of 0.1. Temporal aggregation uses masked mean pooling to handle variable-length sequences.
@@
Temporal structure is preserved through sinusoidal encodings:
@@
where the blade weight (0.677) was tuned via Optuna to balance gradients and label counts.
@@
Augmentation is applied after splitting. Methods include temporal jitter ($\Delta t\sim\{-2,\ldots,2\}$), Gaussian noise ($\epsilon\sim\mathcal{N}(0,0.05^{2})$), rotational noise ($\theta\sim U(-0.05,0.05)$), and scaling ($s\sim U(0.95,1.05)$). Minority moves are oversampled up to 400 per class, and the majority blade class “6” is downsampled to two-thirds of others.
@@
Hyperparameters were tuned with Optuna [1]. Across 300 trials, we explored model size, loss weighting, weight decay, and other settings. The top runs were analyzed for recurring patterns, and their best-performing settings were combined into the final model.
4.5 Dynamic temporal windowing
During training and cross-validation, FERA-MDT is optimized on trimmed single-fencer sequences corresponding to labeled exchanges. At deployment, however, we typically receive longer, untrimmed per-fencer pose tracks extracted from bout footage, and the model must decide both how many frames to group into an action and which labels to assign. Fencing actions vary substantially in duration: half-steps may last 4–6 frames, while lunges or compound blade actions can span 30–40 frames. Using a single static window either truncates long actions or dilutes short ones with irrelevant context. FERA-MDT therefore uses dynamic temporal windowing during inference.
Given a per-fencer pose sequence $\{x_{t}\}_{t=1}^{T}$, where $T$ can range from a short clip to an entire bout segment, we process frames in segments beginning at index $s$, with initial window $w_{0}=1$. For each window [s,s+w1][s,s+w-1], FERA-MDT produces move logits and a blade-line distribution. If the maximum move confidence is below a threshold $\tau$ (per-class thresholds are tuned via validation), the window expands by one frame, $w\leftarrow w+1$, up to $w_{\max}=40$. If no confident prediction emerges by $w_{\max}$, the window is reset to $w_{0}$ and the start index advances by one frame, $s\leftarrow s+1$. For half-steps, the system checks for a subsequent full step before finalizing the label, to avoid over-fragmenting continuous footwork into many very short predictions.
This adaptive strategy allows short actions to be recognized promptly, while longer actions accumulate sufficient context. It reduces boundary errors and improves recognition of compound actions such as step forward + beat, which are often ambiguous when only partial motion is visible. Because $s$ can start at any frame and the window can grow up to $w_{\max}$, the procedure applies to any sequence length $T$: the same classifier trained on trimmed segments acts as a local detector over a continuous pose stream. Predicted move intervals across consecutive windows are merged using non-maximum suppression on overlapping time spans, yielding a final per-fencer sequence of moves and blade positions aligned with the clip or bout timeline.
4.6 FERA-LM: Semantic Reasoning Interface
After generating move and blade predictions, the system requires a mechanism to bridge the gap between semantic labels and logical rule application. FERA-LM serves as this interface. Rather than functioning as a standalone referee, FERA-LM is designed to validate the semantic representational power of the detected actions.
Using DeepSeek-R1-Distill-Qwen-14B as the teacher model [7], we generate referee decisions with a structured prompt that (i) lists the relevant foil rules retrieved via a FAISS index [13], (ii) provides the left and right fencer move sequences with blade positions, and (iii) requests a two-part answer: a single-line decision (Left / Right / None) and a brief explanation referencing the rules and moves. Responses are constrained to be concise and deterministic for consistent outputs [24].
The student model is DistilGPT-2, trained with causal language modeling on the teacher’s outputs [11]. This yields a smaller, cheaper model while retaining decision quality. Both teacher and student are supervised to reproduce the official referee decision derived from the scoreboard.
@@ 5 Results
We evaluate FERA via (i) an end-to-end example showing how the full pipeline produces referee decisions, and (ii) quantitative comparisons of FERA-MDT with baseline models, including ablations.
5.1 End-to-end Logic Consistency
We illustrate one clip where the referee ruled that the left fencer held priority. FERA-MDT correctly identifies the actions for both fencers, and FERA-LM produces a decision and explanation that align with the referee’s judgment.
Crucially, this example demonstrates the logical consistency of the decoupled framework. Even if intermediate move detection errors occur (e.g., missing a minor blade beat), the language model can often recover the correct priority flow if the primary attack sequence (Lunge vs. Retreat) is correctly identified. This suggests that the kinematic features captured by FERA-MDT form a sufficient semantic vocabulary for fencing logic.
@@ 5.1 End-to-end Logic Consistency
(a) Left step forward + beat, right stepping back.
(b) Left lunge, right counterattack.
@@
Explanation: The left fencer initiates the attack with a step forward while the right fencer retreats. The left fencer then performs a lunge, hit (134–146) and the right fencer responds with counterattack, hit (138–148). Under the rules, priority remains with the initial attacker and does not transfer to the right fencer.
5.2 Evaluation metrics
We report four classification metrics and three calibration metrics.
@@ 5.2 Evaluation metrics
  • Macro-F1 $\tfrac{1}{C}\sum_{c}\text{F1}_{c}$: averages F1 over classes, emphasizing rare moves.
  • Weighted-F1 $\sum_{c}\tfrac{n_{c}}{\sum_{j}n_{j}}\,\text{F1}_{c}$: weights each class by frequency.
  • Micro-F1 $\tfrac{2\sum_{c}TP_{c}}{2\sum_{c}TP_{c}+\sum_{c}FP_{c}+\sum_{c}FN_{c}}$: global F1 over all labels.
  • Hamming loss $\tfrac{1}{NC}\sum_{i=1}^{N}\sum_{c=1}^{C}\mathbf{1}[y_{i,c}\neq\hat{y}_{i,c}]$: fraction of misclassified labels; lower is better.
For class $c$,
@@
ECE summarizes average calibration quality; MCE highlights worst-case miscalibration; the Brier score combines calibration and sharpness.
@@ 5.3 Move detection model comparisons
Table 2 compares FERA-MDT with three baselines. FERA-MDT consistently outperforms all baselines on Macro-F1 and Hamming loss, indicating stronger performance on both common and rare moves. BiLSTM comes closest, as recurrent architectures capture short-term motion well, but the encoder-only structure of FERA-MDT appears better suited to the multi-label, long-range dependencies in our setting. The baseline Transformer underperforms slightly despite similar depth, suggesting that pooling and loss weighting matter. TCN achieves the highest Weighted-F1, reflecting a bias toward dominant classes such as step forward and hit, which is undesirable for refereeing, where rare actions also matter.
Table 3 reports calibration metrics. BiLSTM and the baseline Transformer are substantially worse calibrated than FERA-MDT and TCN. Between the latter two, TCN attains lower ECE and MCE but a higher Hamming loss, indicating that its apparent calibration partly reflects overspecialization on frequent classes. FERA-MDT offers a better trade-off between calibration and rare-class recognition.
@@
Overall, FERA-MDT attains the best combination of macro-F1, Hamming loss, and Brier score among the tested models, making it the strongest candidate for downstream referee assistance.
@@ 5.4 Co-occurrence analysis
Referee decisions depend on combinations of actions, so we assess whether models reproduce the ground-truth co-occurrence structure. Figures 58 show co-occurrence mappings. The left matrix in each plot shows ground-truth labels, and the right shows model predictions. FERA-MDT produces mappings that most closely mirror the ground truth; baselines tend to over-activate frequent moves and underrepresent rarer ones.
@@ 5.5 Ablations
We perform ablations on FERA-MDT while keeping all other settings fixed. We report classification (Macro-F1) and calibration (ECE, MCE, Brier) metrics averaged over five folds [8, 3].
Table 4 summarizes the results. The full configuration achieves the highest Macro-F1 (0.549) with strong calibration (ECE = 0.152, MCE = 0.458, Brier = 0.135). Removing per-class thresholds reduces Macro-F1 to 0.511, confirming their importance. Using only the 12 normalized joints (24D) yields Macro-F1 of 0.534, showing that engineered features improve performance but are not strictly required; calibration, however, deteriorates, indicating that features primarily enhance reliability.
Augmentation strategies provide complementary gains. Individually, noise-only, temporal-only, and feature-specific augmentations yield Macro-F1 scores of 0.542, 0.525, and 0.537, respectively. Combining all strategies produces the best robustness and top score (0.549). Figures 10 and 10 illustrate how single augmentations can be unstable across folds, underscoring the need for combined augmentation and calibration.
Per-class temperature scaling improves calibration without changing F1, and down-weighting blade loss slightly benefits both calibration and classification.
Overall, FERA-MDT’s complete design—feature engineering, diverse augmentation, calibrated thresholds, and tailored loss weighting—is necessary to achieve both strong classification and reliable confidence estimates.
@@ 6 Limitations and Future Work
The primary constraint is data. With no public datasets containing extensive move annotations, we relied on manual labeling. Performance improved steadily as the dataset grew: with $\sim$1,000 annotated segments, F1 plateaued near 0.40, while doubling to 2,000 segments raised F1 to 0.549. Long-tail actions remain a bottleneck; expanding the dataset, especially for rare moves, should further improve accuracy.
FERA also depends on generic pose detectors that do not support blade detection. A blade-specific extractor would allow FERA-MDT to incorporate precise blade positions and more reliable contact cues. At the reasoning stage, FERA-LM inherits FERA-MDT’s errors, since teacher generation and student distillation are both conditioned on predicted moves. Improving move detection should therefore also improve language-model decisions.
Evaluating language-based reasoning is another open issue. Our dataset includes final referee decisions (via synchronized scoreboard changes), and both teacher and student are supervised to reproduce these outcomes. However, we cannot disentangle FERA-LM’s contribution from upstream recognition errors, and we lack ground-truth textual justifications from referees. Broadcast footage does not provide access to the referee’s phrasing or intermediate reasoning, so we cannot directly measure explanation quality. The end-to-end example in Section 5 should thus be viewed as a qualitative demonstration. A fuller study would collect structured referee judgments or explanations and evaluate both decision accuracy and explanation usefulness, as in recent VARS work.
Generalization also remains a challenge. Our dataset consists exclusively of professional fencers, often recorded from elevated camera angles, which may limit adaptability to beginners and to different viewpoints. Future work may incorporate 3D joint extraction to mitigate camera-angle sensitivity and extend testing to broader fencer populations.
Finally, we do not use athlete-disjoint splits, as we lack metadata to link clips to individual fencers. While our dataset spans roughly 130 competitors, some overfitting to individuals is still possible.
Overall, these limitations suggest a roadmap: larger and more diverse datasets, explicit blade detection, quantitative evaluation of language-based reasoning, improved distillation from cleaner labels, and greater attention to generalization. Addressing these factors will enable more dependable referee assistance and broader coaching applications.
@@ 7 Conclusion
Modern foil refereeing is entirely manual; only the hit lights are automated. No systems currently assist or validate human decisions in this setting. Prior work on fencing move recognition has been limited to small subsets of footwork actions and often uses specialized sensors.
FERA is a first prototype referee assistant that operates directly from monocular video. We present a pose-based pipeline for recognizing multi-label moves and blade positions for individual fencers, a new dataset with frame-level annotations from real competition footage, and a decoupled system that integrates motion recognition with rule-grounded reasoning. FERA-MDT’s macro-F1 of 0.549 highlights the difficulty of this setting but also demonstrates the feasibility of Transformer-based motion understanding under real-world conditions. By aggregating recognized actions and blade positions from both sides of the piste, and by running the same model over continuous pose streams using dynamic temporal windowing, the same machinery can support both referee assistance and coaching or tactical analysis. We will release the FERA-Dataset annotations, derived pose features, and baseline code to encourage further work on pose-based motion understanding and AI-assisted sports officiating.